A monochromatic light beam of wavelength λ = 550 nm is directed at a film of water at 23° from the vertical as shown. What is the minimum thickness of the film such that I see the brightest reflection?
1
Expert's answer
2020-11-16T07:49:43-0500

As per the given question,

Wavelength of the wave "(\lambda)=550nm"

"\theta=23^\circ"

"\sin \theta =n\frac{\lambda}{d}"

from bright fringe n = 1

Now, substituting the values

"\sin 23^\circ =1\times \frac{550\times 10^{-9}m}{d}"

"\Rightarrow d = \frac{5.5\times 10^{-9}}{0.39}"

"\Rightarrow d = 14.1\times 10^{-9}m"

"d=0.0141\mu m"